# Ego-Exo4D: Understanding Skilled Human Activity from First- and Third-Person Perspectives

### v3

@@ Results
Table 22: Breakdown of results for demonstrator proficiency estimation across scenarios. Top-1 accuracies per scenario for the TimeSFormer model performing best on v1 (K400) and on v2 (EgoVLPv2) versions. All models are trained on the v1 train set.
Val results
Ego Exos Ego + Exos
Method mAP0.25 mAP1.0 Avg. mAP0.25 mAP1.0 Avg. mAP0.25 mAP1.0 Avg.
Random tips/good exec. 0.47/0.48 4.70/5.23 2.21/2.46 0.47/0.48 4.70/5.23 2.21/2.46 0.47/0.48 4.70/5.23 2.21/2.46
Uniform tips 0.50/0.49 5.31/5.28 2.49/2.48 0.50/0.49 5.31/5.28 2.49/2.48 0.50/0.49 5.31/5.28 2.49/2.48
Uniform good exec. 0.45/0.43 4.87/4.79 2.32/2.27 0.45/0.43 4.87/4.79 2.32/2.27 0.45/0.43 4.87/4.79 2.32/2.27
ActionFormer [190] 1.57/0.95 9.60/6.33 5.14/3.27 1.19/1.08 6.99/7.50 3.76/3.84 0.96/0.97 6.72/7.03 3.43/3.57
Test results
Ego Exos Ego + Exos
Method mAP0.25 mAP1.0 Avg. mAP0.25 mAP1.0 Avg. mAP0.25 mAP1.0 Avg.
Random tips/good exec. 0.49/0.45 5.23/4.72 2.47/2.20 0.49/0.45 5.23/4.72 2.47/2.20 0.49/0.45 5.23/4.72 2.47/2.20
Uniform tips 0.44/0.45 5.08/5.18 2.35/2.39 0.44/0.45 5.08/5.18 2.35/2.39 0.44/0.45 5.08/5.18 2.35/2.39
Uniform good exec. 0.45/0.46 4.50/4.62 2.11/2.17 0.45/0.46 4.50/4.62 2.11/2.17 0.45/0.46 4.50/4.62 2.11/2.17
ActionFormer [190] 1.08/1.04 7.20/7.56 3.75/3.87 0.91/1.14 6.08/7.36 3.14/3.87 0.91/1.14 6.77/7.90 3.47/4.04
[…]
Contribution Statement
@@
Santhosh Kumar Ramakrishnan co-led the proficiency estimation benchmark, co-developed the task formulation, and advised the baseline development. Gedas Bertasius co-led the proficiency estimation benchmark, co-developed the task formulation, and advised the baseline development. Arjun Somayazulu developed the demonstrator proficiency estimation baselines. Abrham Gebreselasie developed the demonstration proficiency estimation baselines. Maria Escobar contributed to the task definition and the baseline design. Eugene Byrne contributed to the task definition and the baseline design. Miguel Martin developed an interface for obtaining proficiency estimation scores from the recruited experts. Suyog Jain contributed to an annotation pipeline for demonstration proficiency estimation. Devansh Kukreja built the render flow to generate frame-aligned videos of each camera for each take as model input. Kristen Grauman contributed to the task formulation. Lorenzo Torresani contributed to editing this section.
13.D   Ego Pose
This family of tasks is motivated by recovering the skilled movements of experts in the extreme setting of monocular ego-video input.
13.D.1 3D human body pose from ego-video
Annotations
In the automatic ground truth generation phase, we use off-the-shelf models [25] to predict the 2D bounding boxes from each of the exocentric views. Since there could be multiple people in the scene and we only want to consider the one wearing the egocentric camera, we project the 3D headset location from the MPS output to select which box corresponds to the camera wearer. Then, we run an off-the-shelf 2D human keypoint detector [25] for each bounding box to obtain the 2D keypoints. Finally, we run 3D triangulation with RANSAC to minimize the reprojection errors to obtain the 3D keypoints for the camera wearer.
In the manual annotation phase, we import the undistorted frames and the reprojected 2D keypoints into our multi-view annotation interface.
How good is the auto GT? Between manual and automatic annotations, the body and hand MPJPEs are 3.33 cm and 1.87 cm, respectively, much smaller than the best baseline methods. It is important to note that Ego-Exo4D tackles real-world scenarios with five or fewer cameras rather than controlled environments. This introduces challenges like increased occlusions from body and objects along with limited view of hands from distant cameras. Despite this, our auto generation pipeline surpasses baselines, showcasing robustness and efficacy. Experiments below further show performance boosts across baselines when using automatic ground truth, demonstrating its effectiveness. Note that automatic GT and manual GT are not mutually exclusive, and people can choose whether/how automatic GT is used for training.
Formal task definition
In our ego-pose task, the goal is to estimate the 3D human pose using either an RGB video input sequence $\mathcal{V}_{ego}$, an IMU sensor sequence $\mathcal{T}_{imu}$, or both. For a current frame $t$, given either an egocentric video $\mathcal{V}_{ego}=\{\mathcal{V}^{k},...,\mathcal{V}^{t-1}\}$ or an IMU sequence $\mathcal{T}_{imu}=\{\mathcal{T}^{k},...,\mathcal{T}^{t-1}\}$, where $k$ is a time window in the past, the goal is to predict the human pose $P$ at time $t$, where $P^{t}\in\mathcal{R}^{17\times 3}$ for the 17 annotated joints. Note that at test time, we only estimate the error across the visible annotated joints at frame $t$. Note that, since the EgoPose benchmark is aimed at promoting the development of methods that perform body pose estimation solely from first-person raw video or IMU data, the input excludes egocentric modalities that would unfairly simplify the task (e.g., audio captured from a wearable camera, eye gaze), as well as exocentric video or any signals that can be extracted from it.
Metrics
Baselines
We evaluate several baselines methods. Kinpoly [98] is a physics-based simulator that estimates pose using hard physics constraints applied to an underlying artificial humanoid model. EgoEgo [80] is a diffusion based method which maps a sequence of 6 DoF head estimates to a full sequence of 3D body joint estimates. We also develop a method that combines current Location-based body pose estimation models Bodiffusion [18] and AvatarPoser [64]. Moreover, to gauge the performance of deep-learning-based methods, we implemented a more straightforward approach. We create a static pose baseline, which consists of fixing the 3D human body pose prediction to be the average pose in the training set and translating it according to the IMU sensor. Thus, the fixed prediction matches the camera location at each frame.
  • Kinpoly. Kinpoly [98] proposes to use a simulated humanoid to track head pose and create full-body motion based on action types. Based on the input head-pose and action type (out of push, step, avoid, sit, and locomotion), Kinpoly synthesizes realistic human pose and human-object interactions inside a physics simulator. Different from kinematic-based methods that directly output joint angles or positions for pose estimation, Kinpoly outputs joint torques as the final product and controls a simulated humanoid for pose estimation. The laws of physics and human dynamics regulates the output pose to remove artifacts such as jitter and penetration.
  • EgoEgo. EgoEgo [80] uses a two-step approach for egocentric body pose estimation, by estimating the head pose from the egocentric video first, and then using a diffusion model to generate the full body motion sequence based on the head pose sequence. For head pose estimation, it obtains the initial head pose trajectory using DROID-SLAM [160], and then uses learning-based methods to correct the head pose, including a GravityNet to estimate the additional rotation and a HeadNet with optical flow features as input to estimate the scaling factor to the trajectory. The full body pose is generated with a modified version of DDPM [53] that is conditioned on head pose and trained on AMASS [99]. We show the evaluation of the conditional diffusion part here.
  • Location-based. This baseline is inspired by state-of-the-art methods that use transformer-based models for body pose estimation from sparse inputs[18, 64]. We adapt these methods to utilize 3D positions as opposed to the traditional parametric body model. During the training phase, the model was subjected to $40,000$ iterations, using the Adam optimizer with a learning rate of $1e^{-4}$. The window size for temporal analysis was set at 40 frames, and we minimized the Mean Squared Error (MSE) loss between predicted poses and ground truths. As for the input, our model receives a sequence of head poses captured by the device.
Results
Table 24 shows the evaluation results of all the baseline approaches. First, note that the static pose baseline obtains a significantly higher MPJPE than all the other approaches. This finding suggests that the poses across different scenarios in the dataset are extremely diverse. Thus, attempting to have the same static pose for all test cases is unfeasible. In contrast, the proposed baseline implementations achieve notable enhancements in performance. Table 25 shows the performance of each method per scenario.
While these developments are promising, we believe that further refinement is possible, especially in lower body pose estimation and to ensure temporal consistency in predictions.
Method Validation Test
MPJPE MPJVE MPJPE MPJVE
Static pose 163.94/254.29 - 150.95/215.87 -
EgoEgo [80] 24.35/24.53 0.71/0.78 28.78/26.38 0.57/0.66
Kinpoly [98] 22.66/21.66 0.74/0.86 25.80/24.36 0.60/0.65
Location-based [18, 64] 20.86/20.73 1.87/0.74 19.88/18.51 1.70/0.64
Table 24: Results for the 3D human body pose benchmark. We report the Mean Per Joint Position Error in cm and the Mean Per Joint Velocity Error in m/s for all the baseline approaches on the v1/v2 data releases. Models are trained on the v1 training set.
Scenario EgoEgo Kinpoly Location-based
Basketball 21.36 24.98 19.89
Soccer 23.08 19.09 16.62
Bike repair 30.18 25.19 20.61
Cooking 23.71 20.80 12.65
Health 32.57 29.23 11.63
Dance 20.93 18.03 21.15
Music 33.81 30.30 15.00
Table 25: Body pose estimation Test results per scenario. We report the Mean Per Joint Position Error in cm for the v2 test set.
13.D.2 3D human hand pose from ego-video
Annotations
As for the case of body pose annotation, the 3D human hand pose annotation process also consists of two stages, i.e., the automatic ground truth generation and the manual multi-view keypoint annotation. Compared to the body pose, the main difference in automatic ground truth generation is that we also detect hand keypoints from the egocentric frame, and we use the result from the whole body pose estimation to infer the hand locations when there are multiple people in the scene. Similarly, for manual annotation, besides the exocentric frames, we also show the annotators the egocentric frames to allow them to annotate/correct hand keypoints.
For each annotated joint in manual annotations, we provide the number of views used for triangulation as the indicator of the confidence for the provided ground truth data. Meanwhile, the correction the annotators make for hand joints on ego images can serve as the indicator to understand the difficulty for hand reconstruction from the given ego view.
Formal task definition
The ego hand pose task entails predicting the three-dimensional coordinates of the camera wearer’s hands from the egocentric frames. Frames from the ego view are extracted and undistorted for both training and evaluation. For baselines requiring 2D hand bounding boxes, we project 3D hand joints in camera coordinates to 2D image planes using the provided intrinsic matrix for the crops.
Note that the inputs of the task do not include depth maps, any additional views from ego or exo cameras, camera pose information, and IMU or active range sensor measurements. We explicitly exclude such information for this benchmark to promote the applicability of the methods to general hand pose estimation problems using monocular images.
Metrics
To cope with methods estimating wrist-origin and camera-origin hand pose, the ego hand pose baselines are evaluated according to both the MPJPE and the PA-MPJPE metrics. The MPJPE is camera-relative while the PA-MPJPE calculates the average 3D joint distance after performing Procrustes Alignment on wrist-origin hand poses. Both metrics are reported in millimeter (mm) unit.
Baselines
We implemented and trained several baseline models for ego hand pose estimation. To estimate the 3D hand joint from monocular 2D ego view images, 2D heatmaps can be explicitly estimated and lifted to 3D space, or 3D joints can be directly estimated from extracted 2D features. The feature extractor backbone could be CNN-based or transformer-based. The proposed baseline methods cover different choices of model designs. All the baseline models work on single frame images without temporal information. The baseline models are trained on manual or manual+automatic annotations, and are only evaluated on manual annotations.
  • THOR-net. THOR-net [1] uses Keypoint-RCNN as the feature extractor to obtain 2D information and derive 2D hand keypoints heatmaps explicitly. The method then lifts 2D estimates to the 3D space using GraFormer [197], which is a model consisting of Graph Convolutional layers and Attention layers. We use only the 2D-to-3D pose GraFormer branch in THOR-net to adapt the method to our task. The training takes around 4 hours for the manual dataset on a GeForce RTX 4090 Graphics Card, and around 10 hours for the dataset combining manual and automatic annotations.
  • HandOccNet. HandOccNet [112] uses a ResNet50[51]-based FPN [88] to extract 2D features. The method then uses two Transformer-based modules: Feature Injecting Transformer (FIT) to inject hand information into occluded region, and Self-Enhancing Transformer (SET) to further refine the 2D features. The method proposes a regressor based architecture to produce 2D keypoints, MANO [138] pose, and MANO shape parameters to predict joints and vertices. To accommodate our baseline, only 2D keypoints and 3D joints location losses are used in the training phase. The training takes around 2 hours for the manual dataset on 8 NVIDIA V100 Graphics Cards.
  • POTTER. POTTER [200] proposes Pooling Attention Transformer (PAT) to extract 2D visual features, which significantly reduces the memory and computational cost without sacrificing performances. The method then applies a mesh regression head HybrIK [79] to generate 3D joint and mesh results. The training takes around 43 minutes for manual dataset, and around 4 hours for manual+auto dataset on a GeForce RTX 4090 Graphics Card.
  • METRO. METRO [84] extracts a CNN-based global image features. The method then uses a transformer encoder to jointly model vertex-vertex and vertex-joint interactions, and outputs 3D joint coordinates and mesh vertices simultaneously. Since the training of METRO strongly depends on hand mesh supervision, which is not present in the annotations, we borrowed the checkpoint trained on FreiHand [72] dataset and run the inference only, without training it on our benchmark.
Manual Manual+Auto
MPJPE PA-MPJPE MPJPE PA-MPJPE
METRO* [84] - 20.61 - 20.61
THOR-net [1] 52.05 17.72 49.86 16.00
HandOccNet [112] - 19.74 - 19.47
POTTER [200] 31.73 11.53 29.53 11.16
THOR-net[1] HandOccNet[112] POTTER[200]
Params (M) 59.5 37.22 14.5
MACs (G) 123.6 15.5 5.2
Table 27: Number of parameters and MACS for the different ego hand pose baselines.
Results
We report the MPJPE and PA-MPJPE of the baseline models in Table 26, and their corresponding parameter numbers and multiply-accumulate operations (MACs) in Table 27. We further analyze the error distribution across different hand joints. Figure 43 shows that the thumb finger and finger tips tend to have larger errors, mostl likely because they are occluded or invisible more often.
# visible views 3 4 5 6
PA-MPJPE (mm) 14.01 12.15 11.03 10.02
Table 28: PA-MPJPE for joints that are visible in different number of views (including ego and exo views). Results generated from POTTER [200] evaluation.
For each annotated joint, the manual annotations keep record of the number of views where the joint is visible. The visible 2D observation is then used for triangulation in 3D ground truth generation. This can be taken as an indicator of the uncertainty of the ground truth, and the difficulty level for the estimation of the joint (usually, a joint visible by fewer views indicates that it is more entangled with objects or other part of the hand). Table 28 shows that the PA-MPJPE decreases as the visible number of views increases. To guarantee the ground truth accuracy, all experiments are performed only on joints at least visible in 3 cameras.
Figure 43: Average PA-MPJPE for each joint. Results generated from POTTER [200] evaluation.
Contribution Statement
Kris Kitani co-led the ego-pose benchmark, and provided directional guidance on the task definition, the annotation methodology, and the baseline development. Jianbo Shi co-led the ego-pose benchmark, and provided directional guidance on automatic 3D hand pose generation and development of ego hand pose baseline methods. Maria Escobar led the ego-pose body baseline development, implemented the IMU-based baseline and contributed to experiment analysis. Cristhian Forigua developed the static pose baseline and contributed to the implementation of the IMU-based baseline. Fu-Jen Chu developed the multi-view annotation UI, the hand pose annotation guidelines, and the data preprocessing code for ego-pose annotation; he also trained and evaluated the HandOccNet baseline. Rawal Khirodkar developed the multi-view triangulation and 3D body keypoint estimation pipeline. Zhengyi Luo contributed to the Kinpoly baseline and to the coordinate transform for Aria head poses. Shan Su led the ego-pose hands baseline development, and contributed to automatic 3D hand pose generation, task definition, and annotation development; she also evaluated the baseline using METRO. Suyog Jain developed the annotation pipeline to scale the annotation collection, worked on training the annotators and managed the overall annotation process. Miguel Martin contributed to the automatic ground truth generation pipeline and provided high-level coordination of the body and hands automatic ground truth generation. Jinxu Zhang developed automatic ground truth generation for 3D hand pose; he also trained and evaluated baseline model POTTER. Yiming Huang trained and evaluated the baseline model THOR-net. Zhifan Zhu developed the METRO hand pose baseline method. Jing Huang led the automatic ground truth generation effort, refined body pose annotation guidelines, coordinated ego-pose body baseline development, and ran the EgoEgo body pose baseline. Lorenzo Torresani contributed to editing this section.

### v4

@@ E.5 Proficiency estimation
Table 24: Breakdown of results for demonstrator proficiency estimation across scenarios. Top-1 accuracies per scenario for the TimeSFormer model.
E.6 Ego Pose
CVPR 2024 challenges
We organized two Ego Pose challenges as part of the EgoVis workshop at CVPR 2024, aiming to encourage researchers to explore and utilize the dataset. We had eight participants in total, divided evenly between the Body Pose challenge and the Hand Pose challenge.
Body Pose
There were a total of four submissions to the hand pose challenge, with three outperforming the best baseline method, as shown in Table 25. The best participant outperformed the baseline by 3.19 cm in MPJPE.
MPJPE MPJVE
Location-based (baseline) 18.51 0.64
Levelwise att ViT 18.09 0.62
UCB Ego 17.19 0.57
Multi-Scale Model Fusion 15.32 0.55
Table 25: Leaderboard of Ego-Exo4D Body Pose challenge at EgoVis CVPR 2024.
First place: Multi-Scale model fusion (Baoqi Pei, Yifei Huang, Guo Chen, Jilan Xu, Yicheng Liu, Yuping He, Kanghua Pan, Tong Lu, Limin Wang, Yali Wang, Yu Qiao)
Instead of implementing a new baseline model, this approach leverages the existing location-based baseline and focused on improving performance by fusing predictions from multiple models with varying numbers of transformer layers (ranging from 1 to 6). This approach allowed them to address variations in data distributions and stabilize predictions. By combining the outputs from the different versions of the baseline, they achieved an improved MPJPE of 15.32.
Second place: UCB Ego (Brent Yi, Vickie Ye, Georgios Pavlakos, Lea Müller, Maya Zheng, Yi Ma, Jitendra Malik, Angjoo Kanazawa)
The submission by UCB EGO to the Ego-Exo4D Body Pose Challenge centers around their development of a conditional human motion diffusion model that operates using 30Hz SLAM pose data to directly sample SMPL human body parameters. The model was trained exclusively on the AMASS dataset, which poses a limitation of domain transfer to the Ego-Exo4D dataset. To address this limitation, the team introduced an AdapterNet to estimate floor height and route procedural tasks to the baseline model while focusing the diffusion model on physical activities. This hybrid approach allowed the team to achieve a MPJPE of 17.19 cm.
Third place: Levelwise attention ViT (Congsheng Xu, Jinfan Liu, Yifan Liu, Shuwen Wu)
The authors proposed a model that combines two Vision Transformer (ViT) structures to leverage both coarse-grained and fine-grained information for 3D human pose estimation. They used a baseline ViT model with 8 attention heads and 3 layers for coarse-grained estimation, and a Huge-ViT model with 16 attention heads and 32 layers for fine-grained estimation. They combine the outputs from both ViTs through a weighting strategy. By assigning a higher weight to the fine-grained results, the method balances local detail and global information. This weighting approach resulted in a MPJPE of 18.09 cm.
MPJPE PA-MPJPE
POTTER (baseline) 28.94 11.07
PCIE EgoHandPose 25.51 8.49
Hand3D 30.52 9.30
POTTER-ensemble 28.68 10.24
Table 26: Leaderboard of Ego-Exo4D Hand Pose challenge at EgoVis CVPR 2024.
Hand Pose
There are in total 4 submissions to the hand pose challenge, with 3 outperformed the baseline method POTTER (Zheng et al., 2023) as shown in Table 26. The best participant outperformed the baseline model by 11.85% in MPJPE, and by 23.31% in PA-MPJPE.
First place: PCIE EgoHandPose (Chen et al., 2024)
The authors propose the Hand Pose Vision Transformer (HP-ViT). The HP-ViT comprises a ViT backbone and transformer head to estimate joint position in 3D, utilizing MPJPE and RLE loss function. To be more specific, the model employed regression loss based on RLE (Li et al., 2021) to minimize the gap between output and input distributions. The experiments show that the model with ViT-Huge (Dosovitskiy, 2020) as backbone achieves the best performance. The ensembling model with different setting also contributes to decrease the overall error.
Second place: Hand3D (Pavlakos et al., 2024)
The authors apply recently introduced HaMeR out-of-the-box on the images of the EgoExo4D Challenge, and observe very strong performance. HaMeR is a feed-forward model that takes as input a single image of a hand and hand side (left or right), and estimates a 3D reconstruction of the hand in the form of the MANO parametric hand model (Romero et al., 2017). HaMeR adopts a fully transformerized architecture design using a ViT-H (Dosovitskiy, 2020) backbone, followed by a transformer head. The authors further show the limit of the method, particularly when the occlusions/truncations are very extreme (only a few visible fingers), the wrist location or hand orientation is ambiguous, and finally, if the original hand bounding box crop is not very accurate.
Third place: POTTER-ensemble (Baoqi Pei, Yifei Huang, Guo Chen, Jilan Xu, Yicheng Liu, Yuping He, Kanghua Pan, Tong Lu, Limin Wang, Yali Wang, Yu Qiao)
The authors handle the problems through a multi-scale model fusion method. To be specific, the authors enhance the baseline method POTTER (Zheng et al., 2023) by (1) including multiple upsampling dimensions, specifically 128, 256, and 512, (2) integrating a dynamic pooling operation before the 3D convolution operation of the model, and (3) combining/ensembling these models to improve generalization capabilities.
Appendix F Detailed contribution statement
This project is the result of a large collaboration between many institutions over the last two years. Initial authors represent the leadership team of the project. Kristen Grauman initiated the project, served as the technical lead, initiated the recognition and proficiency benchmarks and expert commentary, and coordinated their working groups. Andrew Westbury served as the program manager and operations lead for all aspects of the project. Lorenzo Torresani led development of the capture domains, initiated the relation and ego-pose benchmarks, and coordinated their working groups. Kris Kitani led development of the multi-camera rig and supported the Ego-Exo4D engineering team on all aspects of the data annotation and organization. Jitendra Malik served as a scientific advisor. Authors with stars () were key drivers of implementation, collection, and/or annotation development throughout the project. Authors with daggers () are faculty and senior researcher PIs for the project.
Camera rig
Aria
Jing Dong and Vijay Baiyya were responsible for obtaining camera poses, calibration, pointclouds and eye gaze using Aria MPS, created the 3D/4D visualizations for the paper and supplementary material, and acted as main contact points from the Aria team throughout the program; with Jing leading the algorithm development and verification, and Vijay leading the Aria MPS workflow and infrastructure development. Jakob Engel acted as technical and scientific advisor, and led the team that built the Aria Localization and Point Cloud algorithms. Kiran Somasundram helped design the capture setup and time-synchronization. Xiaqing Pan helped to align the Aria engineering team to support the EgoExo4D project. Mingfei Yan, Prince Gupta, and Sach Lakhavani acted as product managers of Aria and organizational leads for the successful use of Aria in the program. Kelly Forbes helped setting up agreements and working through the legal requirements of using Aria devices for recording the EgoExo4D dataset across the globe. Richard Newcombe initiated the Aria/Ego4D collaboration and acted as a scientific advisor throughout the program. Furthermore, we want to acknowledge the contribution of the entire Project Aria team as listed in (Engel et al., 2023), including Carl Ren and Sean Diener leading the Aria software and hardware engineering organization, and Renzo De Nardi as technical lead for the Aria device.
Data collection
Carnegie Mellon University Rawal Khirodkar developed the automatic 3D body keypoints extraction pipeline and collected a subset of the soccer, bike mechanic, and cooking sequences for the CMU portion of the dataset. Sean Crane was in charge of the data collection, IRB documents, capturing data, working with participants and processing the data for CMU. Abrham Gebreselasie ran the actionformer-based baseline for demonstration proficiency benchmark. Eugene Byrne served as the engineering lead for the initial design and implementation of the dataset, camera rig, processing pipeline and keystep annotations while at Meta. Subsequently at CMU, he assisted in the recognition benchmarks, implemented Ego-Exo transfer  Li et al. (2021) (1 of the 3 baseline methods for keystep recognition) and the initial implementation of keystep action detection  Zhang et al. (2022), and assisted in annotation/data quality generally.
Indiana University David Crandall oversaw the overall effort at Indiana University, including protocol design and data collection. Weslie Khoo led the IRB protocol design and compliance, arranged logistics such as ordering equipment, and designed and oversaw the cooking scenario data collection. Yuchen Wang and Ziwei Zhao co-led the participant recruitment and data collection for all three scenarios. Ziwei Zhao led data preparation and transfer. We also acknowledge Manasi Swaminathan who assisted with data collection and video synchronization.
Language annotations
Benchmarks
Ego-exo correspondence Manolis Savva co-led the correspondence benchmark and contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Effrosyni Mavroudi co-led the correspondence benchmark and contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Lorenzo Torresani contributed to the task formulation. Sanjay Haresh developed the spatial baselines and contributed to data analysis, experimental results, and paper writing. Yongsen Mao developed the spatiotemporal baselines and contributed to data analysis, experimental results, and paper writing. Suyog Jain formulated the annotation pipeline, developed annotation tools and contributed to annotation guidelines and paper writing. Santhosh Ramakrishnan contributed to the annotation guidelines and the formulation of the annotation pipeline. Xitong Yang contributed to the annotation guidelines and the task definition. We would like to acknowledge Hanxiao Jiang for helpful discussions and preliminary ideas on baseline implementation. Devansh Kukreja built the render flow to generate frame-aligned videos of each camera for each take as model input.
Ego-exo translation Lorenzo Torresani co-led the translation benchmark, developed the task formulation, and advised the baseline development. Judy Hoffman co-led the translation benchmark and advised the baseline development. Feng Cheng led the baseline development and implemented the pix2pix and DiT models for track prediction and clip generation. Mi Luo implemented the GNT baseline and the evaluation pipeline for ego track prediction. Ziwei Zhao contributed the pix2pix baseline for multi-frame input and led the evaluation for ego clip generation. Huiyu Wang advised the baseline development and contributed to the task definition, the baseline design, and the metric selection and analysis.
Fine-grained keystep recognition Tushar Nagarajan co-led the keystep recognition benchmark, co-developed the task formulation, and advised the baseline development. David Crandall co-led the keystep recognition benchmark and contributed to the task formulation. Yale Song co-led the keystep recognition benchmark and led the keystep annotation effort, including design of annotation guidelines, taxonomy development and coordination of annotation workflows; he also contributed to the task formulation, advised baseline design, and facilitated the delivery of EgoVLPv2 pretrained backbone. Triantafyllos Afouras contributed to the taxonomy definition, managed the labeling effort, and developed software for post-processing the annotations. Zihui Xue led the baseline development effort, implemented the TimeSFormer, EgoVLP, VI Encoder and Viewpoint distillation baselines, and performed analysis of results. Eugene Byrne contributed to the taxonomy development and to the Ego-Exo Transfer baseline implementation and analysis. Avijit Dasgupta contributed to the annotation and taxonomy development, and to the early stage of baseline design. Miguel Martin contributed to the annotation and the taxonomy development. Shraman Pramanick contributed the EgoVLPv2 pretrained backbone. Yifei Huang contributed to the early stages of task definition and baseline design. Devansh Kukreja built the render flow to generate frame-aligned videos of each camera for each take as model input, and to produce video collages for annotations. Kristen Grauman contributed to the task formulation.
Energy-efficient multimodal keystep recognition Tushar Nagarajan led the energy-efficient multimodal benchmark, co-developed the task formulation, and advised the baseline development. Sagnik Majumder led the baseline development effort and contributed all baseline implementations and analysis of results for the benchmark. Merey Ramazanova developed the energy profiler used to evaluate all baselines and contributed to the experimental analysis. Mitesh Kumar Singh helped design the energy formula. Miao Liu and Shengxin Cindy Zha initiated this benchmark and developed an early version of the task formulation.
Procedure understanding Antonino Furnari led the procedure understanding benchmark, and contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Giovanni Maria Farinella contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Luigi Seminara contributed to the annotation guidelines, the baseline design, and paper writing; he also developed tools for data annotation, and the baselines for the benchmark. Francesco Ragusa contributed to the annotation guidelines, the baseline design, and paper writing; he also developed tools for data annotation, and the baselines for the benchmark. Kumar Ashutosh contributed to the annotation guidelines, baseline design, and development of data annotation tools. Michael Wray contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Siddhant Bansal contributed to the task definition, the annotation guidelines, the baseline design, and paper writing. Gene Byrne contributed to the task definition, the annotation guidelines and the baseline design. Tushar Nagarajan contributed to the task definition, and the baseline design.
@@ Appendix F Detailed contribution statement
Ego-exo proficiency estimation Santhosh Kumar Ramakrishnan co-led the proficiency estimation benchmark, co-developed the task formulation, and advised the baseline development. Gedas Bertasius co-led the proficiency estimation benchmark, co-developed the task formulation, and advised the baseline development. Arjun Somayazulu developed the demonstrator proficiency estimation baselines. Abrham Gebreselasie developed the demonstration proficiency estimation baselines. Maria Escobar contributed to the task definition and the baseline design. Eugene Byrne contributed to the task definition and the baseline design. Miguel Martin developed an interface for obtaining proficiency estimation scores from the recruited experts. Suyog Jain contributed to an annotation pipeline for demonstration proficiency estimation. Devansh Kukreja built the render flow to generate frame-aligned videos of each camera for each take as model input. Kristen Grauman contributed to the task formulation.
Ego pose Kris Kitani co-led the ego-pose benchmark, and provided directional guidance on the task definition, the annotation methodology, and the baseline development. Jianbo Shi co-led the ego-pose benchmark, and provided directional guidance on automatic 3D hand pose generation and development of ego hand pose baseline methods. Maria Escobar led the ego-pose body baseline development, implemented the IMU-based baseline and contributed to experiment analysis. Cristhian Forigua developed the static pose baseline and contributed to the implementation of the IMU-based baseline. Fu-Jen Chu developed the multi-view annotation UI, the hand pose annotation guidelines, and the data preprocessing code for ego-pose annotation; he also trained and evaluated the HandOccNet baseline. Rawal Khirodkar developed the multi-view triangulation and 3D body keypoint estimation pipeline. Zhengyi Luo contributed to the Kinpoly baseline and to the coordinate transform for Aria head poses. Shan Su led the ego-pose hands baseline development, and contributed to automatic 3D hand pose generation, task definition, and annotation development; she also evaluated the baseline using METRO. Suyog Jain developed the annotation pipeline to scale the annotation collection, worked on training the annotators and managed the overall annotation process. Miguel Martin contributed to the automatic ground truth generation pipeline and provided high-level coordination of the body and hands automatic ground truth generation. Jinxu Zhang developed automatic ground truth generation for 3D hand pose; he also trained and evaluated baseline model POTTER. Yiming Huang trained and evaluated the baseline model THOR-net. Zhifan Zhu developed the METRO hand pose baseline method. Jing Huang led the automatic ground truth generation effort, refined body pose annotation guidelines, coordinated ego-pose body baseline development, and ran the EgoEgo body pose baseline.